Operation Summary for September 16, 2011 - September 19, 2011 |
The Integrated Luminosity for the period from 9/12/11 to 9/19/11 was 27.36 pb-1 (inverse picobarns). NuMI reported receiving 2.6E18 protons on target during this same period. |
Friday September 16 The day shift began with the Tevatron (TEV) conducting a 450 GeV 3x3 study store, with Operations monitoring the stash (358.8E10), with Pbar stacking (68E10), with the Main Injector (MI), Linac, and Booster operating normally, with NuMI off for a target change, with the MiniBooNE experiment taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using study stores. During the day shift, Operations terminated the 3x3 study store; this was the last of the TEV low energy study stores. Operators prepared D0 and CDF for 30-minute accesses at 9:12 AM. Operations established beam to the Meson FTBF experiment T-922 at 9:40 AM. Operators put the TEV into a dry squeeze at 10:44 AM and through a wet squeeze at 11:22 AM. Operations put the TEV into HEP shot setup at 12:35 PM and established store 9125 at 1:40 PM with an average initial luminosity (AIL) of 320.63E30. The I- Source vacuum valve closed at 1:49 PM; it opened. During the evening shift, Operations disabled beam to T-922 at 6:01 PM. At 9:13 PM, a Pbar Debuncher kicker stopped pulsing; operators investigated and then called in a Kicker person who swapped out the high voltage supply and made other adjustments. Pbar resumed stacking at 11:01 PM. Operators reported short downtimes due to a Linac RF station (LRF3), a Linac quadrupole magnet power supply (L:QPS115), a Linac Klystron RF station (KRF5), and a Booster RF station (BRF4), Saturday September 17 The midnight shift began with Operations monitoring store 9125 (91.22E30) and the stash (360.2E10), with Pbar stacking (8.3E10), and with MiniBooNE taking beam. During the midnight shift, Operations terminated store 9125 at 1:26 AM, put the TEV into shot setup, and established store 9126 at 2:48 AM with an AIL of 354.02E30. Operations established beam to T-922 at 5:06 AM. During the day shift, Operations disabled beam to T-922 at 9:30 AM. Operations terminated store 9126 at 9:49 AM; operators prepared D0 for an access. At 10:04 AM, MiniBooNE suffered a horn strip line air pressure trip for the first of two times in eight minutes; operators notified a MiniBooNE person who reported that a sensor switch had gone bad. Operators bypassed the sensor; the MiniBooNE person will come in and investigate further. Water Works personnel reported at 12:44 PM that they were regenerating column #2 at CUB. At 2:22 PM, an operator and a MiniBooNE person accessed the MI-12 service building and unstuck the horn air pressure switch. D0 came out of access at 2:33 PM. Operations put the TEV through a dry squeeze at 3:16 PM and a wet squeeze at 3:57 PM. During the evening shift, Operations put the TEV through another wet squeeze at 5:17 PM. Operations put the TEV into shot setup at 5:38 PM and established store 9131 at 6:45 PM with an AIL of 365.8E30. At 9:01 PM, store 9131 aborted; there was no quench. Operators investigated and reported that the TEV sector C3 dump beam permit had tripped off due to the loss of a 5-volt power supply; operators called in an EE Support person who had the dump repaired by 11:54 PM. Sunday September 18 The midnight shift began with Operations monitoring the stash, with Pbar stacking, and with MiniBooNE taking beam. During the midnight shift, Operations put the TEV into shot setup at 2:18 AM and established store 9134 at 3:18 AM with an AIL of 207.21E30. During the day shift, operators monitored the accelerator complex. During the evening shift, Operations terminated store 9134 at 6:06 PM, put the TEV into shot setup, and established store 9135 at 7:22 PM with an AIL of 374.8E30. Operators reported at 8:19 PM that the Pelletron suffered the first of two full discharges in 13 minutes; operators contacted a Pelletron person and then made some small adjustments. At 11:26 PM, LRF1 tripped off; operators reset the station, but then reported a banging noise coming from the power amplifier transmission line. Operators called in Linac personnel and put Pbar into standby. Monday September 19 The midnight shift began with Operations monitoring store 9135 and the stash, with Pbar and the MiniBooNE experiment off due to Linac being down. During the midnight shift, Linac personnel reported that they had replaced a fuse for the LRF1 driver anode; beam returned at 1:40 AM. At 2:29 AM, the TEV suffered a seven-house quench; operators investigated and believed the cause of the quench was from a kicker prefire. Cryo System personnel began the TEV sector C4 cold compressor pump down at 3:36 AM and then continued on with the rest of the houses. Operators put the TEV separators in conditioning mode, per a TEV person’s instruction. At 5:27 AM, the I- Source vacuum valve closed; it reset and opened. Operators reported at 7:48 AM that all of the TEV cryo systems were cold. The Plans for Monday The plans for today are to stack, stash, and store. The Run Coordinator will allow TEL2 studies, MI 8 GeV studies, and Booster studies. |
|
More Information Comments and Suggestions |
| last modified 09/19/2011 email Fermilab |
| Security, Privacy, Legal | |